1)

Find the EAR in each of the following cases. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16. Use 365 days in a year.) Number of Times Compounded Stated Rate (APR) 11 .25% 14.75 17.25 13.25 Effective Rate (EAR) Quarterly Monthly Daily Semiannually2)

3)
Answer #1

1)

Calculation of EAR when Stated rate =-11 .25% Number of compounding periodn4 (Compounded quarterly) So Effective Rate (EAR) r
Answer #2



SOLUTION :


1.


Stated rate                 Number of times                   Effective rate

  (APR)                          compounded                            (EAR)



  11.25%                     Quarterly                       (1 + 0.1125/4)^4 - 1 = 0.1173 = 11.73%


 14.75%                      Monthly                      (1 + 0.1475/12)^12 - 1 = 0.1579 = 15.79%


 17.25%                      Daily                       (1 + 0.1725/365)^365 - 1 = 0.1882 = 18.82%


 13.25%                      Semiannually                  (1 + 0.1325/2)^2 - 1 = 0.1396 = 13.69% 



2.



Stated rate                                                     Number of times                 Effective rate

  (APR)                                                              compounded                          (EAR)



 (( 1 + 0.1450)^(1/4) - 1)*400       = 13.77%           Quarterly                        14.50%

 

(( 1 + 0.1050)^(1/12) - 1)*1200    = 10.03%           Monthly                          10.50%


(( 1 + 0.1050)^(1/52) - 1)*5200    =   9.99%           Weekly                           10.50%


((1 + 0.0850)^(1/365) - 1)*36500 =   8.16%            Daily                                8.50% 



3.


Let interest rate be r in decimals.


So,


PV = A *((1+r)^n - 1)/(r(1+r)^n) 

=> 760 = 76*((1+r)^32 - 1)/(r(1+r)^32) 

By trial and error , r = 0.0944 = 9.44%


Interest rate is 9.44% (ANSWER)

answered by: Tulsiram Garg